OpenStudy (anonymous):

Given f(x) = 3x - 1 and g(x)= -x + 6, find f(-2) + g(5).

OpenStudy (anonymous):

lol, you are in the totally wrong section but lucky you i was looking at this one. for f(-2) you replace x with -2 in the problem that is f(x). so 3 * -2 - 1 = -6 - 1 = -7. now we have to do the g one. go g(5) = -5+6 = -1. but we aren't done yet!!! so f(-2) + g(5) = -7 + -1 = -7 -1 = -8